A flexible loop has a radius of 15.3 cm and it is in a magnetic field of B = 0.176 T. The loop is grasped at points A and B and stretched until its area is zero. It takes 0.237 s to close the loop. (a) What is the magnetic flux through the loop before it is streched? (b) What is the magnetic flux through the loop after it is stretched? (c) What is the magnitude of the average induced electromotive force ϵ in the loop during the stretching process? Problem 2. (0.1 pt). A tightly wound circular coil has 59 turns, each of radius 0.235 m. A uniform magnetic
Answer #1

a) Magnetic Flux = BA = 0.176 × (π × 0.153 × 0.153) = 1.29 × 10-2

Wb.

b) Magnetic flux through the loop after it is stretched = 0 Wb

c) E = dB/dt = 1.29 × 10-2/0.237 = 5.44 × 10-2 V.